OpenStudy (anonymous):

How do you find all the roots (four) to this equation knowing that one of them is 1+3i? x^4+3x^3+6x^2+2x-60 How do you find all the roots (four) to this equation knowing that one of them is 1+3i? x^4+3x^3+6x^2+2x-60 @Mathematics

OpenStudy (anonymous):

1-3i is another. divide them out (after multiplying together) using polynomial long division or synthetic division and then factor. Alternately, graph the function and find the other two that way.

OpenStudy (anonymous):

Okay, what do you mean divide them out after multiplying them together?

OpenStudy (turingtest):

(1-3i)(1+3i)=1-3=-2 so x+2 is a factor can you do the synthetic division?

OpenStudy (anonymous):

yes, i did the synthetic division for -2 and the bottom level ended up being 1 -5 16 -30 0

OpenStudy (anonymous):

where do I go from there to find the last root?